OpenStudy (anonymous):

Solve. x – 6 = –14 A.–20 B.20 C.8 D.–8

OpenStudy (anonymous):

@mathmath333 @Nnesha @Jamie182

OpenStudy (anonymous):

I will give you a metal

OpenStudy (johnweldon1993):

Hint...if you add 6 to both sides of that equation...

OpenStudy (anonymous):

D

OpenStudy (anonymous):

is it a or d

OpenStudy (anonymous):

x-6=-14 Transpose 6 to the other side then solve.

OpenStudy (johnweldon1993):

Or think of it this way "Some number, minus 6....gives you -14" what is that number?

OpenStudy (anonymous):

hey @MadisonEmery need help? ^_^

OpenStudy (anonymous):

solving for x???

OpenStudy (anonymous):

yes @Captain_America_Army

OpenStudy (anonymous):

Add 6 to both sides. x – 6 = –14 x - 6+6 = -14+6 x = -8

OpenStudy (anonymous):

so.. @YesThisIsDog was correct ^_^

OpenStudy (anonymous):

thx can you help me with more

Nnesha (nnesha):

too much enabling wouldn't be good for someone future

OpenStudy (anonymous):

Btw this is @just_one_last_goodbye ^_^ im sure u tagged me for help... and im really sorry i didn't go to the other question u tagged me for

OpenStudy (anonymous):

it's fine can you help me with more @Captain_America_Army

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Solve. 5 = –(–z + 3) A.–8 B.8 C.–2 D.2

OpenStudy (anonymous):

@Captain_America_Army

OpenStudy (anonymous):

–(–z + 3) = 5 Refine: z - 3 = 5 Add 3 to both sides: z - 3 + 3=5+ 3 z = 8

OpenStudy (anonymous):

What is the solution to the equation?

OpenStudy (anonymous):

z = 8

OpenStudy (anonymous):

What is the solution to the equation? 5+(x-12)=18 X=

OpenStudy (anonymous):

5+(x-12)=18 Refine: x - 7 = 18 Add 7 to both sides: x - 7 + 7 = 18 + 7 x = 25

OpenStudy (anonymous):

Serena parachutes from a plane at a height of 20,000 ft above sea level. Several minutes later she lands on the ground 15 ft below sea level. What was the total vertical distance that Serena parachuted? how many feet

OpenStudy (anonymous):

all you have to do... is add the 2 numbers... 20,000 + 15 = _____ ft <---(answer)

OpenStudy (anonymous):

20015

OpenStudy (anonymous):

The variable t represents the temperature at 3 p.m. Which equation would best model this word problem? The temperature at the summit of Mt. Mitchell dropped 23°F between 3 p.m. and 9 p.m. The temperature at 9 p.m. was –10°F. What was the temperature at 3 p.m.? A. t – 23 = –10 B.10 – 23 = t C.10 + 23 = t D.23 – t = –10

OpenStudy (anonymous):

on the feet one is the anwser 20015

OpenStudy (anonymous):

20,015

OpenStudy (anonymous):

^ correct

OpenStudy (anonymous):

The answer for this one is.... 23 +(-10) = 13 so look for something similar to this expression they like to trick people by rearranging the numbers ^_^ so it's not C or D

OpenStudy (anonymous):

so d

OpenStudy (anonymous):

or a which one @Captain_America_Army

OpenStudy (anonymous):

it was d

OpenStudy (anonymous):

sorry ;-; was doing something